OpenStudy (anonymous):

Find the value of k such that L(3,-1) M(0,5) N(k,-4)

OpenStudy (anonymous):

Complete the question !

OpenStudy (anonymous):

sorry, Find the value of k such that L(3,-1) M(0,5) N(k,-4) are collinear

OpenStudy (anonymous):

The points L , M and N are collinear if : \[\Large \overrightarrow {LM}=\alpha\Large \overrightarrow {LN}\] So : \[\Large\begin{pmatrix}0-3\\5-(-1)\end{pmatrix}=\alpha\begin{pmatrix}k-3\\-4-(-1)\end{pmatrix}\] So : \[\Large\begin{cases}\alpha(k-3)=-3\\-3\alpha=6\end{cases}\] From the 2nd equation we get : \[\Large \alpha=-2\] And from the 1st we get : \[\Large -2(k-3)=-3\] so : \[\Large k-3=\frac32\] so : \[\Large k=\frac32+3=\frac 92\]

OpenStudy (anonymous):

i don't know how to use alpha

OpenStudy (anonymous):

substitute ... alpha ....

OpenStudy (anonymous):

From the 2nd equation we get the value of alpha, then we replace the value of alpha in the 1st equation

OpenStudy (anonymous):

ok, thanks.